A hot-air balloon is rising straight up with a speed of 3.15 m/s. A ballast bag is released from rest relative to the balloon when it is 6.89 m above the ground. How much time elapses before the ballast bag hits the ground?
Answer #1

0=3.15 - gt

t= 3.15/9.8 =0.32 sec

0 = (3.15)^2 - 2*9.8*x

x=0.5m

hight = 6.89 + (0.5) = 7.4 m

7.4 = (1/2)*9.8*t^2

t=1.23 s

net time = 1.23 + 0.32 =1.55 sec